find steady state temperature of black body disc having 2cm and mass 7g at 45 degree C SURROUNDING temperature, when kept initially at 28 degree C .hence determine Stefan's constant


1
Expert's answer
2020-12-15T11:54:36-0500

"By\ Boltzman's\ Law \\

\\
\frac{\partial Q _{out} }{\partial t} =( \ A \Pi^{2} )\sigma (T^{4}_{sum_{}}-T^{4})\\
Gain\ rate\\
\frac{\partial Q _{in} }{\partial t} = \frac{\partial }{\partial t }mc\Delta T =mc\frac{\partial T}{\partial t}\\
Equating\ both \\
\frac{\partial Q _{out} }{\partial t} = \frac{\partial Q _{in} }{\partial t}\\
\frac{\partial T}{\partial t} =\frac{A\pi ^{2}}{mc}\sigma T_{sum}^{4}-\frac{A\pi ^{2}}{mc}\sigma T^{4}\\

\frac{\partial T}{\partial t} + \frac{A\pi ^{2}}{mc}\sigma T^{4} =\frac{A\pi ^{2}}{mc}\sigma T_{sum}^{4}\\
T (t) = \frac{1}{\sqrt[3]{\frac{3A\pi ^{2}} {mc}\sigma t +(\frac{1}{T-T_{sum}})3}} +T_{sum}\\
In\ steady -state\\
t \to \infty\\
T=T_{sum}=45^{\circ}C\\
on \ 318K\ (Steady)"